# Lite-Mind: Towards Efficient and Versatile Brain Representation Network

### 06 Dec 2023 v1

@@ 4.3 Retrieval pipeline
Test set retrieval. The cosine similarities between voxel embeddings and image embeddings are directly calculated.
LAION-5B retrieval. Image embeddings are generated based on voxel CLS token by Diffusion Projector for online retrieval. Because our DFT backbone uses the same one-way contrast loss as CLIP, we consider that the DFT backbone outputs voxel embeddings similar to text embeddings, which we consider as voxel-level text. We found that when searching on the large-scale dataset LAION-5B, img2img performs better than text2img and can find more similar images. Therefore, we used a similar method to represent the disjointed voxels of the backbone output, using a DALLE · 2[29] similar diffusion model to obtain the image representation and retrieve it on LAION-5B.
Zero-shot classification. Image retrieval is performed on novel classes in the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
4.4 Contrastive Learning
Due to the inherent differences between different modalities, we use CLIP’s contrastive loss to train the DFT backbone. To perform img2img retrieval on LAION-5B, we use MSE loss to constrain the generation of approximate image representations. Sampling a batch $\mathcal{B}$ from voxel-image pairs, the contrastive loss is defined below:
$L_{contr}=-\frac{1}{|\mathcal{B}|}\sum_{s=1}^{|\mathcal{B}|}\log\frac{\exp(f_{s}^{\top}\cdot\mathcal{V}_{s}/{\tau})}{\sum_{i=1}^{|\mathcal{B}|}\exp({f_{s}^{\top}\cdot\mathcal{V}_{i}}/{\tau})},$ (5)
where $f_{s}$ is the $s$-th voxel representation, $\mathcal{V}_{s}$ is its corresponding image representation, $\tau$ is a temperature factor.
Accordingly, the training objective is defined below:
$L=\alpha L_{contr}+\beta L_{mse}$ (6)
When training LAION-5B retrieval tasks, $\beta$ is a non-zero value; When training test set retrieval, $\beta=0$. Due to the lightweight performance of the model, we trained two separate models for fMRI-image retrieval and image-fMRI retrieval, corresponding to two types of one-way losses.
@@ 5.1 Dataset
Figure 3: Partial search results of Lite-Mind on all 982 test images for Subject 1. Even without diffusion prior and heavy MLP backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates. The number below each image represents the similarity score.
@@
Ozcelik… [25] 257 separate linear regression models 3B 21.1% 30.3%
MindEye [33] MLP Backbone 940M 13.3% 63.1%
@@
MindEye [33] MLP Backbone+Prior 3B 93.4% 90.1%
Lite-Mind(ours) DFT Backbone 12.5M 94.3% 97.1%
(a) Quantitative comparison of Lite-Mind retrieval performance against other models for Subject 1. Image retrieval refers to the percent of the time the correct image was retrieved out of 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval. Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities(see Appendix B for remaining subject models).
@@
5.2 Implementation details
To fairly compare different backbones, we did not use any data augmentation methods, such as voxel mixing loss MixCo and image slicing enhancement. Due to MindEye only disclosing the MLP Backbone performance of Subject 1, only the performance of Subject 1 is provided in Table 0(a). The experimental results of more subjects are shown in Appendix B. All of our experiments were conducted on a single Tesla V100 32GB GPU. More experimental details and hyperparameter settings can be found in Appendix A.
6 Results
6.1 fMRI/image retrieval
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings in the training set. If a paired image embedding is retrieved, the retrieval is considered correct. fMRI retrieval is the opposite process mentioned above. Note that there are many semantically and visually similar images in the NSD dataset test set, which are considered to be similar in CLIP space. Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image. For test retrieval, we adhered to the identical methodology as Lin et al. [19] and MindEye [33] to compute the retrieval metrics presented in Table 0(a). For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images. The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set. We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples. The experimental results are shown in Table 0(a). We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone. The detailed results and discussions are in Appendix B.
As shown in Table 0(a), compared to MLP Backbone+Projector our DFT Backbone improves retrieval accuracy by 5.5% and 12.2% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value. Furthermore, our DFT Backbone retrieval performance even surpasses MindEye’s MLP Backbone+Prior without any image enhancement methods, proving that frequency domain calculations better capture global features.
@@ 6 Results
Figure 4: Partial search results of Lite-Mind on LAION-5B retrieval for Subject 1. The left column marked by a red box in every two columns represents the original image seen by the subject, and the right column represents the image retrieved by LAION-5B.
@@
Table 1: Zero-shot visual stimulus classification on GOD dataset. The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%). The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from [7]. MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B for retrieval performance on the GOD dataset).
6.2 LAION-5B retrieval
Because Lite-Mind can also align image representation space, we can expand the retrieval task to larger datasets, such as LAION-5B [32]. For LIAON-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS token of CLIP. The final layer CLIP ViT-L/14 embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2]. For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS token, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP). Unlike MindEye, in the LAION-5B retrieval process, we used the diffusion model of DALLE·2 prior to conducting img2img retrieval.
To compare the ability of LAION-5B retrieval to replace image reconstruction, we compared the performance indicators of reconstruction with previous image reconstruction methods. The experimental results are shown in Table 0(b). The specific evaluation indicators in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstructions; SSIM is structural similarity index metric [39]; EfficientNetB1(”Eff”) [37] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
Experiments show that our diffusion projector performs better than the projector of MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has a parameter size of 8M (MindEye-CLS backbone still has 125M). The results indicate that efficient DFT backbones can also assist downstream retrieval tasks. Partial search results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval performance and results of each subject are shown in Appendix B.
6.3 GOD zero-shot classification
Since the collective prime values and data volume of fMRI data are much smaller than those of the NSD dataset in most cases, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability. The voxel values collected by participants in the GOD dataset range from 4133 to 4643, much smaller than that of the NSD dataset, and the training set has only 1200 fMRI-image pairs(4.8% of NSD training set).
We conduct zero-shot classification by using fMRI-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification). We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset. We also replicated the MLP backbone and projector methods of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction in voxel vector length, and MindEye still has a parameter quantity of 210M. As shown in Table 1, the experimental results indicate that our proposed Lite-Mind still guarantees versatility and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training. On the contrary, due to the sudden decrease in data size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data volume and data augmentation for voxel value-wise mappings in heavy MLP Backbone. More retrieval and classification results are provided in Appendix B.
@@ 6.4 Ablations and visualization
Architectural Analysis. We assess the effectiveness of different modules within Lite-Mind to investigate the lightweight and efficient performance in this session. All ablation experiments are for Subject 1 on the NSD dataset.
@@
We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of DFT on retrieval efficiency(Table 2). We even completely removed the DFT layer and aligned the patches directly to the last hidden layer of CLIP ViT-L/14 with a tiny MLP, still achieving a forward retrieval accuracy of 65.8%, indicating that patches have more concentrated information than voxels. As the number of DFT layers gradually deepens, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual. We interestingly found that the accuracy of brain retrieval tends to converge faster than the image and only 6 layers of DFT are needed to approach the highest value. We have also conducted more ablation experiments on data volume to demonstrate the robustness of Lite-Mind. Refer to Appendix B for more ablations.
@@
Visualization. We visualize the diffusion model representation of LAION-5B using T-SNE in Figure 5. The graph shows that the diffusion model successfully transformed the voxel representation learned from contrastive learning into an image representation, and the two representations were well fused together for img2img retrieval on the LAION-5B dataset. We have also conducted more visualization experiments in Appendix C to demonstrate the efficiency of DFT Backbone.
DFT depth Parameters Image Retrieval Brain Retrieval
Pyramid(18) 12.5M 94.3% 97.4%
18 9.9M 91.8% 97.3%
12 6.6M 91.0% 97.2%
6 3.4M 90.6% 97.2%
1 0.7M 83.5% 94.0%
0 0.2M 65.8% 74.7%
Table 2: Ablation experiments on Discrete Fourier Transform (DFT) layers revealed a positive correlation between the depth of the DFT layer and enhanced forward retrieval accuracy. Simultaneously, augmenting the filter pyramid of the receptive field demonstrated an additional improvement in retrieval accuracy, provided there is a judicious increase in the parameter count.
@@
Figure 5: T-SNE of representations of 982 test fMRI voxels by diffusion projector on LAION-5B retrieval for Subject 5. The diffusion model plays a role in bringing vectors closer, enabling img2img retrieval on the LAION-5B dataset.
@@ 7 Conclusion
In this work, we propose Lite-Mind, an extremely lightweight brain representation extraction network based on Fourier transform for cross-modal fMRI-image retrieval. Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli. With high retrieval accuracy, Lite-Mind can transform many downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification. Meanwhile, Lite-Mind’s dependence on data volume and voxel values is less than that of larger models, demonstrating excellent universality across different fMRI datasets.
For future work, we should explore the relationship between different sizes of patches and brain area structures and explore the impact of cortices other than the visual cortex, such as ventral, on retrieval accuracy. In addition, versatile cross-subject models should be studied to efficiently project fMRI representations of different subjects into the same joint space or transfer from subject models with larger sample sizes to models with sparse sample sizes. At the same time, how to use an efficient model to collaboratively train multiple fMRI datasets (across subjects) is also a promising direction that can be explored in the future.

### 12 Mar 2024 v2

@@ 3.3.4 Downstream tasks
Test set retrieval. The cosine similarities between voxel embeddings from DFT Backbone and image embeddings from CLIP ViT are directly calculated.
LAION-5B retrieval. Voxel CLS embeddings from the DFT backbone are translated to image CLS embeddings through the Diffusion Projector for online LAION-5B retrieval.
Zero-shot classification. Image retrieval is performed on novel classes on the test set, and the similarities between the retrieved images and simple CLIP class text prompts are calculated for classification tasks.
Figure 3: Partial retrieval results of Lite-Mind on all 982 test images for Subject 1. With 12.5M DFT Backbone, Lite-Mind can still find the exact Top-1 image pair from the test set of 982 images with 89.3% accuracy and 95.5% accuracy of image-fMRI retrieval(random chance = 0.1%) and can distinguish among confusable candidates. The number below each image represents the similarity score. See more cases include failure retrieval in Appendix C.4.
@@ 4.1 Dataset
4.2 Implementation details
To fairly compare different backbones, we did not use any data augmentation methods, such as voxel mixing loss MixCo and image slicing enhancement. Due to MindEye only disclosing the MLP Backbone performance of Subject 1, only the performance of Subject 1 is provided in Table 0(a). The experimental results of more subjects are shown in Appendix B.1. All of our experiments were conducted on a single Tesla V100 32GB GPU. More experimental details and hyperparameter settings can be found in Appendix A.
5 Results
5.1 fMRI/image retrieval
Image retrieval refers to retrieving the image embeddings with the highest cosine similarity based on voxel embeddings on the test set. If a paired image embedding is retrieved, the retrieval is considered correct. fMRI retrieval is the opposite process mentioned above. Note that there are many semantically and visually similar images on the NSD dataset test set, which are considered to be similar in CLIP space. Whether the model can correctly retrieve as MindEye tests the fine-grained alignment ability to the image. For test retrieval, we adhered to the identical methodology as Lin et al. [21] and MindEye [35] to compute the retrieval metrics presented in Table 0(a). For each test sample, we randomly selected 299 images from the remaining 981 images in the test set and calculated the cosine similarity between the voxel embeddings and 300 images. The retrieval accuracy refers to the proportion of successful retrieval of corresponding images in the 982 voxel embeddings of the test set. We adjusted the random number seed of 30 randomly selected images to average the accuracy of all samples. The experimental results are shown in Table 0(a). We also conducted retrieval experiments on the remaining three subjects to demonstrate the universality of DFT Backbone. The detailed results and discussions are in Appendix B.2.
As shown in Table 0(a), compared to MLP Backbone, our DFT Backbone improves retrieval accuracy by 5% and 14.9% for two retrieval ways, indicating that the fine-grained representation of the image comes from the rich representation of the last hidden layer of CLIP rather than MLP’s excessive attention to each voxel value. Note that our DFT Backbone is closest to MindEye’s MLP Backbone with only contrastive learning, while Prior in MLP Backbone+Prior includes MSE loss for image reconstruction.
@@ 5 Results
Ozcelik… [27] 257 separate linear regression models 3B 29.9% 21.4%
MindEye [35] MLP Backbone 940M 89.6% 82.2%
@@
MindEye [35] MLP Backbone+Prior 1.2B 93.4% 90.1%
Lite-Mind(ours) DFT Backbone 12.5M 94.6% 97.1%
(a) Quantitative comparison of Lite-Mind retrieval performance against other models for Subject 1. Image retrieval refers to the hit rate of correct retrieval from 300 candidates, given the associated brain sample (chance=0.3%); vice-versa for brain retrieval. Lite-Mind only uses a parameter quantity of 12.5M, achieving extremely high retrieval performance without using any model to close vectors of different modalities(see Appendix B.1 for remaining subject models).
@@
Table 1: Main retrieval results on the NSD and LAION-5B dataset.
5.2 LAION-5B retrieval
Although the retrieval scale on NSD is large enough, we can still expand it to larger datasets, such as LAION-5B [34]. The final layer CLIP ViT-L/14 CLS embeddings for all 5 billion images are available at https://knn.laion.ai/ and can be queried for K-nearest neighbor lookup via the CLIP Retrieval client [2]. For LAION-5B retrieval, we train another similar Lite-Mind-8M, aligning voxels to the CLS embeddings of CLIP. For each test sample, we conduct a retrieval strategy in the same way as MindEye (first retrieve 16 candidate images using CLS embeddings, and the best image is selected based on having the highest cosine similarity to the fMRI voxel embeddings aligned to the final hidden layer of CLIP). Unlike MindEye, in the LAION-5B retrieval process, we used the Diffusion Projector to conduct image-to-image retrieval.
To compare the performance of large-scale dataset retrieval, we use metrics of image reconstruction for evaluation with previous image reconstruction methods. The experimental results are shown in Table 0(b). The specific evaluation metrics in the table are as follows: PixCorr represents pixel-wise correlation between ground truth and retrieval/reconstruction images; SSIM is structural similarity index metric [40]; EfficientNetB1(”Eff”) [39] and SwAV-ResNet50(”SwAV”) [4] refer to average correlation distance; all other metrics refer to two-way identification (chance = 50%).
Experiments show that our DFT Backbone performs better than MindEye in high-level metrics, as shown in Table 0(b), although our backbone only has 8M parameters(MindEye-CLS backbone still has 125M. Note that both MindEye and Lite-Mind perform LAION-5B retrieval based on CLS alignment, resulting in a smaller number of parameters compared to full models). The results indicate that an efficient DFT backbone can also assist downstream retrieval tasks. Partial retrieval visualization results of Lite-Mind on LAION-5B retrieval for Subject 1 are shown in Figure 4, and the specific retrieval and visualization results of each Subject are shown in Appendix B.2.
5.3 GOD zero-shot classification
Since the collective fMRI data are much fewer and shorter than those of the NSD dataset usually, we conducted zero-shot classification tasks on a smaller fMRI dataset to verify the Lite-Mind’s generalization ability. The voxel lengths collected by Subjects on the GOD dataset range from 4133 to 4643, approximately 30% of NSD, and there are only 1200 fMRI-image pairs on the training set (4.8% of NSD).
We conduct zero-shot classification by using fMRI-to-image retrieval to find the specific image and using simple prompt text templates ”An image of [class]” to obtain the category of retrieved images(CLIP has a classification accuracy of 76.2% on ImageNet without pre-training on it, so it can be seen as zero-shot classification). We train our Lite-Mind-15.5M to align voxels to images on the GOD dataset. We also replicated the MLP backbone and projector of MindEye, changing the residual block of $4096\times 4096$ to $1024\times 1024$ to accommodate the reduction of voxel length(We empirically found that the original $4096\times 4096$ performed even worse), and MindEye still has 210M parameters. As shown in Table 2, the experimental results indicate that our proposed Lite-Mind still guarantees robustness and establishes a new state-of-the-art for zero-shot classification on the GOD dataset without complex text prompt templates or auxiliary text representation training. On the contrary, due to the sudden decrease in dataset size and voxel length, the performance of MindEye significantly decreases, indicating excessive reliance on training data size for voxel value-wise mappings with large MLP Backbone. More retrieval and classification results are in Appendix B.3.
@@ 5.3 GOD zero-shot classification
Figure 4: Partial retrieval results of Lite-Mind from LAION-5B for Subject 1. The left column marked by a red box in every two columns represents the original image seen by the Subject, and the right column represents the image retrieved from LAION-5B.
@@
Table 2: Zero-shot visual stimulus classification on the GOD dataset. The test set contains 50 categories that have no overlapping with the training set (top-1 chance=2.0%). The results for CADA-VAE, MVAE, MMVAE, MoPoE-VAE, and BraVL are taken from [7]. MindEye(211M) is a smaller version consisting of MLP backbone and projectors, with a residual block size of $1024\times 1024$ (See Appendix B.3 for retrieval performance on the GOD dataset).
@@ 5.4 Ablations and visualization
We assess the effectiveness of different modules within Lite-Mind to investigate the lightweight and efficient performance in this session. All experiments below are for Subject 1 on the NSD dataset.
@@
Architectural Analysis. We trained multiple DFT Backbone models with different depths to evaluate the impact of the number of layers of fMRI Spectrum Compression on retrieval efficiency(Table 3). As the Filter Blocks layers gradually deepen, the number of model parameters increases, the image retrieval accuracy gradually improves, and the improvement speed tends to be gradual. We interestingly found that the accuracy of brain retrieval tends to converge faster than the image retrieval and only 6 layers of Filter Blocks are needed to approach the highest value. We have also conducted more ablation experiments on hyper-parameters to demonstrate the robustness of Lite-Mind. Refer to Appendix B.4 for it.
Depth FLOPs(G) Parameters Image Retrieval Brain Retrieval
MindEye 5.66 996M 88.8% 84.9%
1 0.05 0.7M 83.5% 94.0%
6 0.22 3.4M 90.6% 97.2%
12 0.43 6.6M 91.0% 97.2%
21 0.80 12.5M 94.6% 97.4%
Table 3: Ablation experiments on fMRI Spectrum Compression revealed a positive correlation between the depth of the Filter Blocks and enhanced forward retrieval accuracy.
Module Analysis. We conducted ablation experiments on two main modules of DFT Backbone in Table 4. We completely removed the Filter blocks and aligned tokens directly to the image embeddings with fMRI token decoupling, achieving a forward retrieval accuracy of 65.8%, indicating that the importance of frequency domain filtering for denoising and compression. Then we replaced FreMLP with the FC layer in the real domain, where tokens are flatten into a tensor for alignment, similar to the last layer of the MLP backbone of MindEye. The model performance rapidly declined, indicating that tokens processed by DFT Backbone have better mapping performance with FreMLP in the frequency domain.
Module Image Retrieval Brain Retrieval
DFT Backbone 94.6% 97.4%
w/o Filter Blocks 65.8% 74.7%
w/o FreMLP 43.4% 47.8%
Table 4: Ablation experiments on modules of DFT Backbone.
Embedding Dimension.We conducted additional experiments to verify the impact of embedding dimensions on retrieval accuracy, as shown in Table 5. Among them, two different CLIP models (i.e., CLIP ViT/B-32 and CLIP ViT/L-14) were used to extract image embeddings, with a total of four embedding dimensions, by CLS embeddings and the last hidden layer, respectively. From the results, it can be seen that the retrieval accuracy of Lite-Mind is higher when the embedding dimensions of the image are longer and the representation of the image is richer. The results verify that the fine-grained alignment we mentioned above comes from the rich representation of the image, rather than the fMRI voxel value fully connected heavy MLP Backbone.
Embeddings Dimension Parameters Retrieval
[…]
ViT-B/32 CLS 512 6.7M 57.5% 61.8%
ViT-L/14 CLS 768 6.7M 60.3% 64.2%
ViT-B/32 Hidden 50×512 10.4M 91.1% 96.5%
ViT-L/14 Hidden 257×768 12.5M 94.6% 97.4%
Table 5: Retrieval performance with different CLIP embeddings for Subject 1 on the NSD dataset.
Cerebral cortex. In order to explore the impact of different cortical regions on retrieval accuracy, we used Takagi’s [38] method to extract the stream ROI on the NSD dataset, covering the visual cortex of nsdgenal, and trained the individual retrieval DFT Backbones for visual cortices (early, lateral, parietal, ventral). As shown in Table 6, it can be observed that the early visual cortex has the greatest impact on retrieval accuracy. Although with voxel length of 5917, Lite-Mind can still achieve retrieval accuracy of 85.0% and 93.4%, which is consistent with Takagi’s research. They demonstrated that visual stimuli are almost dominated by the early visual cortex and indicate that nsdgenal still has a low signal-to-noise ratio, proving fully connected MLP Backbones are unnecessary. Interestingly, for other cortical regions, both lateral and ventral have some impacts on retrieval accuracy, while parietal shows little.
Region Voxel Length Parameters Retrieval
[…]
early 5917 12.4M 85.0% 93.4%
lateral 7799 12.4M 26.9% 29.0%
parietal 3548 12.3M 9.6% 11.2%
ventral 7604 12.3M 18.1% 22.4%
@@
Visualization. We visualize the CLS embeddings for LAION-5B using T-SNE in Figure 5. The figure shows that the diffusion projector successfully transformed the voxel CLS embeddings learned from contrastive learning into the image CLS embeddings, and the two kinds of embeddingss are well fused together for image-to-image retrieval on the LAION-5B dataset. We have also visualized the alignment process of voxel embeddings by contrastive learning on the NSD dataset in Appendix C.3, and learned filters of Filter Blocks in fMRI Spectrum Compression in Appendix C.5.
@@
Figure 5: T-SNE of embeddings of 982 test fMRI by diffusion projector on LAION-5B retrieval for Subject 1. In the figure, image-clip represents image CLIP CLS embeddings, while voxel-fft and voxel-diffusion represents voxel CLS embeddings by DFT Backbone or Diffusion Projector. The diffusion projector plays a role in bringing vectors closer, enabling image-to-image retrieval on the LAION-5B dataset.
6 Limitations and Future Work
Due to inherent differences between Subjects, Lite-Mind still needs to train Subject-specific models. Although our exploratory experiments have shown that a common model can achieve comparable retrieval accuracy on four Subjects, it’s still necessary to ensure sufficient training data. Further design beyond the model level is still needed for cross-Subject models, to achieve Subject-specific few-shot or even zero-shot. In addition, Lite-Mind performs worse on Subjects with shorter voxel lengths than Subjects with longer, consistent with previous research, which lacks relevant neuroscience theoretical evidence. This provides a promising avenue for the interpretability of deep learning in neuroscience.
@@ 7 Conclusion
In this work, we propose Lite-Mind, an extremely lightweight brain representation learning paradigm based on Fourier Transform for cross-modal fMRI-to-image retrieval. Our DFT Backbone is an efficient means of obtaining fine-grained representation alignment between fMRI signals and visual stimuli. With high retrieval accuracy, Lite-Mind can transform many downstream fMRI tasks into retrieval-based tasks, such as zero-shot classification. Meanwhile, Lite-Mind’s dependence on dataset size and voxel length is less than larger models, demonstrating excellent robustness across different fMRI datasets.